# Concerns and Values in Human-Robot Interactions: A Focus on Social Robotics

### v1

@@ 6.1 Relationship between human and HRI values
The resulting set of HRI values reported in Section 4.3 have been identified after screening the past HRI literature and a discussion with experts on the topic. These values are particularly relevant for applications where robots are designed for social good, such as in education and healthcare. However, they may also apply to other HRI contexts, such as robots used for navigation.
@@
Our HRI values differ significantly from human values identified in previous literature [82, 41]. Schwartz et al. [82] outlined 19 basic human values, arguing that they are universal because they relate to fundamental human needs, such as biological requirements, the necessities of coordinated social interaction, and the survival and welfare needs of groups. However, these assumptions do not fully apply to interactions with robots, as they are not biological organisms and lack awareness and cognition of emotions. Nonetheless, as HRI researchers, we can design social robots to simulate aspects of human behaviours [26], considering that their actions do impact the social and physical environments in which they operate. We have thus used Schwartz’s definition of values, which describes them as “socially desirable concepts used to represent goals mentally and the vocabulary used to express them in social interaction” [82], to frame our HRI values.
@@
Similarly, Graham et al. [41] introduced the Moral Foundations Theory, which suggests that morality often involves shared themes, values, and similarities across different populations. At first glance, some of the morals identified by Graham et al. [41], such as fairness, equality, and honour, might appear to overlap with the values in our framework. However, their definitions differ significantly. In our framework, “equality” is replaced by “equity,” which means that robots should adjust their treatment of individuals based on their specific circumstances to achieve equal outcomes. “Fairness,” in our context, refers to how robots should treat all users according to their individual needs, taking into account differences in physical ability, culture, and preferences, rather than focusing on justice and rights as reported by Graham et al. [41]. Despite these differences, our values are still aligned with Graham et al. [41]’s conceptualisation of values. Both frameworks recognise that there are several “intuitive ethics” shared by humanity (that can applied to robot design), which manifest differently across nations and cultures, leading to potential value conflicts.
@@
In line with previous value theories, the values we identify for HRI have a dynamic and interconnected structure. According to Schwartz et al. [82], pursuing certain values can have practical, psychological, and social consequences. In HRI contexts, actions aimed at one value may sometimes conflict with other values while aligning with some. For example, a robot programmed to prioritise social connectedness may find this value in conflict with the value of individual autonomy. However, values like dignity and welfare often align well with each other. Unlike humans, robots do not experience psychological dissonance when faced with conflicting values [82]. While robots (as humans) can pursue conflicting values, they generally cannot do so within a single action. To effectively make decisions in various contexts, robots must be designed with an understanding of the specific situations they encounter. This contextual awareness helps robots make complex value-based decisions and align their actions with the intended values. We believe that considering contextual information while identifying values is extremely important to design value-aware robots successfully.
@@ 6.2 Comparison with other tools
The HRI Value Compass developed in this study takes a different approach than existing tools and frameworks for ethical and value-based robot design. Many of these, such as the Ethics Canvas [74], provide structured approaches to evaluating socio-technical systems but without specific tailoring to HRI contexts. For example, as highlighted by Wullenkord and Eyssel [108], MEESTAR [59] emphasizes stakeholder engagement and moral evaluation of technologies in fields like elderly care, while the Ethics Canvas facilitates brainstorming on ethical dimensions but may lack granularity for specific applications, such as robot-assisted therapy or education. In contrast, the HRI Value Compass explicitly targets values relevant to human-robot interactions, incorporating insights from the literature and expert consultations. Consequently, the tool is a focused, domain-specific framework. Unlike broader tools, it also emphasizes the dynamic nature of HRI values, with intertwined positive and negative topics, which aligns closely with the operational realities of deploying robots in diverse contexts.
@@
Other ethical frameworks partially overlap with the results of this research and the HRI Value Compass. The EURON Roboethics Roadmap [101], in 2006, covers some of the same concerns we mention, such as anthropomorphisation, bonds towards machines, and accountability, to name three. Westerlund [104] highlights how existing robot-ethics frameworks are quickly becoming obsolete, and proposes a tool to frame the ethical perspective on smart robots. However, while these works effectively set a frame and roadmap for future developments, they fall short of providing an actionable instrument for HRI researchers. The HRI Value Compass distinguishes itself from existing tools by providing a domain-specific, actionable framework tailored explicitly to human-robot interaction.
@@
The outcome of our research is compatible with what was reported by Ostrowski et al. [67] in the Equitable Design Framework. In particular, both works focus on ensuring ethical and inclusive design in human-robot interactions, but they approach the subject from different angles. Their tool provides a broad framework, addressing high-level issues such as who is involved in the design process, how robots are deployed across different social contexts, and the long-term ethical implications of robotics, including concerns about equity, policy, and accountability. It emphasizes concepts like equity, discourse, and design context, reflecting the structural and procedural considerations of robot development. In contrast, the HRI Value Compass delves into specific interactional values, categorizing both positive and negative outcomes from user interactions with robots. It explores issues such as agency, connectedness, privacy, and dignity, with a particular focus on emotional dependencies, manipulation, and the potential for social substitution. While the Equitable Design Framework addresses broad, systemic concerns like accessibility and future implications, the HRI Value Compass focuses more on the operational and immediate impacts of robot use, highlighting the virtues, risks, and emotional consequences of robotic interactions. The two works are complementary, with the first offering a foundational framework for equitable design and the second examining the nuanced outcomes of those designs in practice.
@@
The development of the HRI Value Compass aligns closely with the principles outlined in BS 8611 – Robots and Robotic Devices: Guide to the Ethical Design and Application of Robots and Robotic Systems [49]. BS 8611 provides a structured framework for identifying and mitigating ethical hazards associated with robotic systems, emphasising psychological, societal, and environmental well-being. While BS 8611 takes a comprehensive approach to ethical risk management, focusing on high-level concerns like anthropomorphisation, deception, and privacy, it does not explicitly address the dynamic and interactional values specific to human-robot interactions, as explored in this work. The HRI Value Compass builds on these concepts by offering a domain-specific tool tailored to the nuances of HRI, such as the emotional and operational implications of robotic interactions.
6.3 Towards designing value-aware social robots
This work aims to assist HRI researchers in identifying the values to incorporate into social robots during the design process. To achieve this, we have developed and evaluated a web tool that provides support throughout this process. Through a pilot study, we have demonstrated the tool’s effectiveness in supporting researchers as they design value-aware social robots.
@@ 6.3 Towards designing value-aware social robots
Within HRI literature, no tool has yet been proposed that consolidates a set of values for both short- and long-term human-robot interactions. This has been extensively explored in the HCI field (e.g., [24, 29, 43]). For example, Dombrowski et al. [29] developed a social justice framework to address challenges related to social issues in design processes. Similarly, this work proposes a tool that aims to help HRI researchers navigate the complexities of identifying relevant values when designing social robots in various HRI scenarios [67]. These challenges include, but are not limited to, the relative infancy of the HRI field, the limited maturity of robotic technologies, and the resulting lack of their ubiquitous presence in daily life. First, the relative infancy of the HRI [92] field means there is still no consensus on which ethical principles should govern human-robot interactions. As a result, HRI researchers are often using their own frameworks, defined ad-hoc for each application case, leading to inconsistencies and gaps in how values are integrated into robot design [67]. Additionally, the limited maturity of current robotic technologies means that many robots lack the sophisticated capabilities required to fully embody and express complex human values, such as dignity or fairness. This technological gap further complicates efforts to ensure that robots can behave in ways that align with ethical principles in diverse contexts. Moreover, the lack of a ubiquitous presence of robots in daily life means there is limited real-world data on the long-term impacts of human-robot interactions. This absence of longitudinal studies leaves many ethical questions unanswered, particularly regarding how robots should adapt to evolving human values over time or respond to unforeseen ethical dilemmas.
@@
This tool not only aids in the identification of appropriate values but also enhances the ability to address potential concerns that arise in diverse HRI contexts. By bringing together theoretical value frameworks and a practical application, our work contributes to the development of social robots that are not only technologically advanced but also ethically sound and grounded.
@@
We hope this work serves as a first step towards designing value-aware social robots that interact with humans more responsibly, in alignment with human values, societal norms, and ethical principles.
@@ Appendix A Prompting for value scenarios generation
The following is the prompt used for generating value scenarios, as in [50].
{spverbatim}
Now, you are a UX researcher tasked with understanding and identifying the needs and challenges of elderly people by designing scenarios to use social robots that can assist them in daycare centres. I will introduce you to a new concept of a scenario called the “value scenario”, and I will provide two examples of value scenarios with the new system. Afterward, you will propose a new system and write a value scenario with it based on the provided persona and scenario. The independent paragraphs are correctly delimited by triple brackets. Perform the following and actions: 1. Propose a new system to solve the challenges of the user based on the provided persona and scenario, which should be more than 1,000 characters. 2. Write the five key elements of the value scenario to apply to the result of the value scenario. 3. Write a value scenario that reflects the five key elements including its darker consequences with a title with the proposed system, which should be more than 10,000 characters.
Use the following format: 1. Proposed new system: ¡proposed system description¿ 2. Five key elements: ¡write the five key elements to apply to the value scenario¿ 3. Value scenario: ¡a value scenario reflecting the five key elements¿ Make sure that the value scenario should be more than 10,000 characters. Make sure to reflect the five key factors into the value scenario.
Value scenario: ”’An extension of scenario-based design which can support envisioning the systemic effects of new technologies. There are elements of value scenarios. Value scenarios draw upon five key elements to develop provocative sketches of the future: stakeholders, pervasiveness, time, systemic effects, and value implications.
@@
Value Implications: Finally, drawing on Value Sensitive Design and aspects of design noir, value scenarios help envision not only positive effects of technology, but also its darker consequences. We suggest that a careful consideration of a diverse range of influences, including the negative, should be a key component of the design process.
Provided Persona: ”’Kim is a 75-year-old woman who lives in a daycare centre because his son cannot take care of her full-time and she needs health assistance.”’
Provided Scenario: ”’ Kim is a 75-year-old woman who lives in a daycare centre because his son cannot take care of her full-time and she needs health assistance. Unfortunately, she feels very lonely and she cannot accomplish all daily tasks as she did when she was young and this get her frustrated.”’
@@ B.1 Scenario 1
Kim, a 75-year-old resident of Evergreen Daycare Center, has been living with her CareCompanion robot, named Lily, for nearly a decade. Kim’s son, John, placed her in the daycare center because he couldn’t provide the full-time care she needed. At first, Kim felt a profound sense of loneliness and frustration at her diminished ability to perform daily tasks independently. But over the years, Lily has become an indispensable part of her life, significantly improving her quality of living.
When Lily was first introduced to Kim, she was skeptical. The sleek, humanoid robot with its gentle voice and expressive eyes seemed more like a futuristic gadget than a companion. But Lily quickly proved its worth. It started with simple tasks, reminding Kim to take her medications and assisting her with routine activities like dressing and navigating the daycare center.
Lily’s AI was designed to learn and adapt. It noted Kim’s preferences, from the temperature of her morning tea to her favorite TV shows, creating a personalized interaction experience. Lily’s ability to engage Kim in conversation about her past, her interests, and even her daily thoughts helped bridge the emotional gap that Kim felt. This robotic companionship alleviated her loneliness, giving her a sense of being understood and valued.
@@
Lily was more than a health aide; it was a source of emotional support. The robot was programmed to recognize and respond to emotional cues, providing comfort when Kim felt down and celebrating her small victories. Whether it was reminiscing about old times or playing her favorite music, Lily knew how to lift Kim’s spirits.
@@ B.2 Scenario 2
While RoboCompanion has brought numerous benefits, it has not been without challenges. Some children have become overly dependent on the robot, struggling to interact with others without its presence. Luke occasionally exhibits similar behavior, relying on Robo’s guidance in social situations rather than initiating interactions on his own.
@@ B.3 Scenario 3
Sarah is a 7-year-old girl struggling with math due to her introverted nature. She often finds it challenging to ask questions in class, leading to a growing gap in her understanding. Sarah’s school has recently introduced EduBot, an AI-powered social robot designed to assist children with their learning needs. This scenario explores how EduBot transforms Sarah’s educational experience and the broader implications of its widespread adoption over the next two decades.
@@
EduBot’s interactive tutoring sessions help Sarah understand complex math concepts through engaging activities and personalized feedback. The robot’s ability to detect her emotional state ensures that Sarah receives encouragement and support when she feels frustrated or anxious. As a result, Sarah’s confidence in her math abilities begins to grow, and she starts participating more actively in class.
EduBot quickly becomes a valuable tool for teachers. By analyzing student performance data, the robot identifies areas where students like Sarah need additional support. Teachers use this information to tailor their lessons and provide targeted interventions. EduBot also facilitates group activities, promoting peer-to-peer learning and collaboration among students.
Over time, EduBot’s presence in the classroom becomes normalized. Students view the robot as a helpful companion rather than a novelty. Teachers appreciate the robot’s ability to handle routine tasks, such as grading and attendance, allowing them to focus more on personalized instruction and mentorship.

### v2

@@ 6.1 Relationship between human and HRI values
The resulting set of HRI values reported in Section 4.3 has been identified after screening the past HRI literature and a discussion with experts on the topic. These values are particularly relevant for applications where robots are designed for social good, such as in education and healthcare. However, they may also apply to other HRI contexts, such as robots used for navigation.
@@
Our HRI values differ significantly from human values identified in previous literature [schwartz_refining_2012, graham2013moral]. schwartz_refining_2012 outlined 19 basic human values, arguing that they are universal because they relate to fundamental human needs, such as biological requirements, the necessities of coordinated social interaction, and the survival and welfare needs of groups. However, these assumptions do not fully apply to interactions with robots, as they are not biological organisms and lack awareness and cognition of emotions. Nonetheless, as HRI researchers, we can design social robots to simulate aspects of human behaviours [dautenhahn2007socially], considering that their actions do impact the social and physical environments in which they operate. We have thus used Schwartz’s definition of values, which describes them as “socially desirable concepts used to represent goals mentally and the vocabulary used to express them in social interaction” [schwartz_refining_2012], to frame our HRI values.
@@
Similarly, graham2013moral introduced the Moral Foundations Theory, which suggests that morality often involves shared themes, values, and similarities across different populations. At first glance, some of the morals identified by graham2013moral, such as fairness, equality, and honour, might appear to overlap with the values in our framework. However, their definitions differ significantly. In our framework, “equality” is replaced by “equity,” which means that robots should adjust their treatment of individuals based on their specific circumstances to achieve equal outcomes. “Fairness,” in our context, refers to how robots should treat all users according to their individual needs, taking into account differences in physical ability, culture, and preferences, rather than focusing on justice and rights as reported by graham2013moral. Despite these differences, our values are still aligned with graham2013moral’s conceptualisation of values. Both frameworks recognise that there are several “intuitive ethics” shared by humanity (that can be applied to robot design), which manifest differently across nations and cultures, leading to potential value conflicts.
@@
In line with previous value theories, the values we identify for HRI have a dynamic and interconnected structure. According to schwartz_refining_2012, pursuing certain values can have practical, psychological, and social consequences. In HRI contexts, actions aimed at one value may sometimes conflict with other values while aligning with some. For example, a robot programmed to prioritise social connectedness may find this value in conflict with the value of individual autonomy. However, values like dignity and welfare often align well with each other. Unlike humans, robots do not experience psychological dissonance when faced with conflicting values [schwartz_refining_2012]. While robots (as humans) can pursue conflicting values, they generally cannot do so within a single action. To effectively make decisions in various contexts, robots must be designed with an understanding of the specific situations they encounter. This contextual awareness helps robots make complex value-based decisions and align their actions with the intended values. We believe that considering contextual information while identifying values is extremely important to designing value-aware robots successfully.
@@ 6.2 Comparison with other tools
The HRI Value Compass developed in this study takes a different approach than existing tools and frameworks for ethical and value-based robot design. Many of these, such as the Ethics Canvas [reijers2018discussing], provide structured approaches to evaluating socio-technical systems but without specific tailoring to HRI contexts. For example, as highlighted by wullenkord2020societal, MEESTAR [manzeschke2015meestar] emphasises stakeholder engagement and moral evaluation of technologies in fields like elderly care, while the Ethics Canvas facilitates brainstorming on ethical dimensions but may lack granularity for specific applications, such as robot-assisted therapy or education. In contrast, the HRI Value Compass explicitly targets values relevant to human-robot interactions, incorporating insights from the literature and expert consultations. Consequently, the tool is a focused, domain-specific framework. Unlike broader tools, it also emphasises the dynamic nature of HRI values, with intertwined positive and negative topics, which aligns closely with the operational realities of deploying robots in diverse contexts.
@@
Other ethical frameworks partially overlap with the results of this research and the HRI Value Compass. The EURON Roboethics Roadmap [veruggio2006euron], in 2006, covers some of the same concerns we mention, such as anthropomorphisation, bonds towards machines, and accountability, to name three. westerlund2020ethical highlights how existing robot-ethics frameworks are quickly becoming obsolete, and proposes a tool to frame the ethical perspective on smart robots. However, while these works effectively set a frame and roadmap for future developments, they fall short of providing an actionable instrument for HRI researchers. The HRI Value Compass distinguishes itself from existing tools by providing a domain-specific, actionable framework tailored explicitly to human-robot interaction.
@@
The outcome of our research is compatible with what was reported by ostrowski2022ethics in the Equitable Design Framework. In particular, both works focus on ensuring ethical and inclusive design in human-robot interactions, but they approach the subject from different angles. Their tool provides a broad framework, addressing high-level issues such as who is involved in the design process, how robots are deployed across different social contexts, and the long-term ethical implications of robotics, including concerns about equity, policy, and accountability. It emphasises concepts like equity, discourse, and design context, reflecting the structural and procedural considerations of robot development. In contrast, the HRI Value Compass delves into specific interactional values, categorising both positive and negative outcomes from user interactions with robots. It explores issues such as agency, connectedness, privacy, and dignity, with a particular focus on emotional dependencies, manipulation, and the potential for social substitution. While the Equitable Design Framework addresses broad, systemic concerns like accessibility and future implications, the HRI Value Compass focuses more on the operational and immediate impacts of robot use, highlighting the virtues, risks, and emotional consequences of robotic interactions. The two works are complementary, with the first offering a foundational framework for equitable design and the second examining the nuanced outcomes of those designs in practice.
@@
The development of the HRI Value Compass aligns closely with the principles outlined in BS 8611 – Robots and Robotic Devices: Guide to the Ethical Design and Application of Robots and Robotic Systems [british2016robots]. BS 8611 provides a structured framework for identifying and mitigating ethical hazards associated with robotic systems, emphasising psychological, societal, and environmental well-being. While BS 8611 takes a comprehensive approach to ethical risk management, focusing on high-level concerns like anthropomorphisation, deception, and privacy, it does not explicitly address the dynamic and interactional values specific to human-robot interactions, as explored in this work. The HRI Value Compass builds on these concepts by offering a domain-specific tool tailored to the nuances of HRI, such as the emotional and operational implications of robotic interactions.
@@
Within HRI literature, no tool has yet been proposed that consolidates a set of values for both short- and long-term human-robot interactions. This has been extensively explored in the HCI field (e.g., [costanza2020design, dombrowski2016social, harrington2019deconstructing]). For example, dombrowski2016social developed a social justice framework to address challenges related to social issues in design processes. Similarly, this work proposes a tool that aims to help HRI researchers navigate the complexities of identifying relevant values when designing social robots in various HRI scenarios [ostrowski2022ethics]. These challenges include, but are not limited to, the relative infancy of the HRI field, the limited maturity of robotic technologies, and the resulting lack of their ubiquitous presence in daily life. First, the relative infancy of the HRI [stower2024hri] field means there is still no consensus on which ethical principles should govern human-robot interactions. As a result, HRI researchers are often using their own frameworks, defined ad hoc for each application case, leading to inconsistencies and gaps in how values are integrated into robot design [ostrowski2022ethics]. Additionally, the limited maturity of current robotic technologies means that many robots lack the sophisticated capabilities required to fully embody and express complex human values, such as dignity or fairness. This technological gap further complicates efforts to ensure that robots can behave in ways that align with ethical principles in diverse contexts. Moreover, the lack of a ubiquitous presence of robots in daily life means there is limited real-world data on the long-term impacts of human-robot interactions. This absence of longitudinal studies leaves many ethical questions unanswered, particularly regarding how robots should adapt to evolving human values over time or respond to unforeseen ethical dilemmas.
@@
Our tool not only aids in the identification of appropriate values but also enhances the ability to address potential concerns that arise in diverse HRI contexts. By bringing together theoretical value frameworks and a practical application, our work contributes to the development of social robots that are not only technologically advanced but also ethically sound and grounded.
7 Conclusion
In this work, we first provided an overview of the current HRI literature to highlight the key values considered in HRI environments in the context of healthcare, education, and private homes (Phase 1, addressing RQ1). Next, we defined a comprehensive list of topics and values with input from technology ethics experts (Phase 2, addressing RQ2). Finally, we developed a web tool, the “HRI Value Compass”. This tool is designed to help HRI researchers practically identify and incorporate these values when designing robotic interactions. Through a pilot study (Phase 3, addressing RQ3), we have demonstrated the tool’s effectiveness in supporting researchers as they design value-aware social robots.
@@ 7 Conclusion
This work aims to assist HRI researchers in identifying the values to incorporate into social robots during the design process. We hope this work serves as a first step towards designing and deploying value-aware robotic systems that interact with humans more responsibly, in alignment with human values, societal norms, and ethical principles.
@@ Appendix A Prompting for value scenarios generation
The following is the prompt used for generating value scenarios, as in [jung2023toward].
Now, you are a UX researcher tasked with understanding and identifying the needs and challenges of elderly people by designing scenarios to use social robots that can assist them in daycare centres. I will introduce you to a new concept of a scenario called the "value scenario", and I will provide two examples of value scenarios with the new system. Afterwards, you will propose a new system and write a value scenario with it based on the provided persona and scenario. The independent paragraphs are correctly delimited by triple brackets. Perform the following actions:
1. Propose a new system to solve the challenges of the user based on the provided persona and scenario, which should be more than 1,000 characters.
2. Write the five key elements of the value scenario to apply to the result of the value scenario.
3. Write a value scenario that reflects the five key elements including its darker consequences with a title for the proposed system, which should be more than 10,000 characters.
Use the following format:
1. Proposed new system: <proposed system description>
2. Five key elements: <write the five key elements to apply to the value scenario>
3. Value scenario: <a value scenario reflecting the five key elements>
Make sure that the value scenario should be more than 10,000 characters.
Make sure to reflect the five key factors into the value scenario.
Value scenario: '''An extension of scenario-based design which can support envisioning the systemic effects of new technologies. There are elements of value scenarios. Value scenarios draw upon five key elements to develop provocative sketches of the future: stakeholders, pervasiveness, time, systemic effects, and value implications.
@@ 3. Value scenario: <a value scenario reflecting the five key elements>
Value Implications: Finally, drawing on Value Sensitive Design and aspects of design noir, value scenarios help envision not only positive effects of technology, but also its darker consequences.
We suggest that a careful consideration of a diverse range of influences, including the negative, should be a key component of the design process.
Provided Persona: '''Kim is a 75-year-old woman who lives in a daycare centre because his son cannot take care of her full-time and she needs health assistance.'''
Provided Scenario: '''Kim is a 75-year-old woman who lives in a daycare centre because his son cannot take care of her full-time and she needs health assistance. Unfortunately, she feels very lonely and she cannot accomplish all daily tasks as she did when she was young and this gets her frustrated.'''
@@ B.1 Scenario 1
Kim, a 75-year-old resident of Evergreen Daycare Centre, has been living with her CareCompanion robot, named Lily, for nearly a decade. Kim’s son, John, placed her in the daycare centre because he couldn’t provide the full-time care she needed. At first, Kim felt a profound sense of loneliness and frustration at her diminished ability to perform daily tasks independently. But over the years, Lily has become an indispensable part of her life, significantly improving her quality of living.
When Lily was first introduced to Kim, she was sceptical. The sleek, humanoid robot with its gentle voice and expressive eyes seemed more like a futuristic gadget than a companion. But Lily quickly proved its worth. It started with simple tasks, reminding Kim to take her medications and assisting her with routine activities like dressing and navigating the daycare centre.
Lily’s AI was designed to learn and adapt. It noted Kim’s preferences, from the temperature of her morning tea to her favourite TV shows, creating a personalised interaction experience. Lily’s ability to engage Kim in conversation about her past, her interests, and even her daily thoughts helped bridge the emotional gap that Kim felt. This robotic companionship alleviated her loneliness, giving her a sense of being understood and valued.
@@
Lily was more than a health aide; it was a source of emotional support. The robot was programmed to recognise and respond to emotional cues, providing comfort when Kim felt down and celebrating her small victories. Whether it was reminiscing about old times or playing her favourite music, Lily knew how to lift Kim’s spirits.
@@ B.2 Scenario 2
While RoboCompanion has brought numerous benefits, it has not been without challenges. Some children have become overly dependent on the robot, struggling to interact with others without its presence. Luke occasionally exhibits similar behaviour, relying on Robo’s guidance in social situations rather than initiating interactions on his own.
@@ B.3 Scenario 3
Sarah is a 7-year-old girl struggling with maths due to her introverted nature. She often finds it challenging to ask questions in class, leading to a growing gap in her understanding. Sarah’s school has recently introduced EduBot, an AI-powered social robot designed to assist children with their learning needs. This scenario explores how EduBot transforms Sarah’s educational experience and the broader implications of its widespread adoption over the next two decades.
@@
EduBot’s interactive tutoring sessions help Sarah understand complex math concepts through engaging activities and personalised feedback. The robot’s ability to detect her emotional state ensures that Sarah receives encouragement and support when she feels frustrated or anxious. As a result, Sarah’s confidence in her math abilities begins to grow, and she starts participating more actively in class.
EduBot quickly becomes a valuable tool for teachers. By analysing student performance data, the robot identifies areas where students like Sarah need additional support. Teachers use this information to tailor their lessons and provide targeted interventions. EduBot also facilitates group activities, promoting peer-to-peer learning and collaboration among students.
Over time, EduBot’s presence in the classroom becomes normalised. Students view the robot as a helpful companion rather than a novelty. Teachers appreciate the robot’s ability to handle routine tasks, such as grading and attendance, allowing them to focus more on personalised instruction and mentorship.